OpenStudy (aryana_maria2323):

Please help! Will FAN AND MEDAL! Find the x-coordinate where the graph of the function f(x) = e(–sinx) has a slope of 0. a. pie b. pie/4 c. pie/2 d. none of these

OpenStudy (aryana_maria2323):

can you help @mathmale

OpenStudy (mathmale):

Hello! Would you mind explaining your " f(x) = e(–sinx) " in words? Is that -sin x meant to be an exponent/power of the base e? If so, please use ^ for exponentiation.

OpenStudy (photon336):

If you've done some calculus, well that's the approach I see I'm wondering if we could just take the derivative of this and then set it equal to zero. \[\frac{ d }{ dx }e^{-sinx} = e^{u}*\frac{ du }{ dx } = e^{-sinx}*\-cos(x) |_{0} \] since we have a product of 2 functions we don't need to guess very much \[\frac{ d }{ dx } e^{-sin(x)}*-Cos(x)\]

OpenStudy (photon336):

I think you could also plot this graph on wolfram alpha and look for the point where the slope is zero.

OpenStudy (mathmale):

Personally I'd rather that aryana find the x-value at which the graph of the given function has a horiz. tan. line (that is, a line with slope 0. wolframalpha would then make for a nice means of checking your answer. So support the idea of finding the derivative of the given function, setting it = 0, and solving for x. Aryana?

OpenStudy (photon336):

that's interesting, yeah it's been a while for me if there is a simpler way to do it then it's better to just use that way.

OpenStudy (mathmale):

Aryana: check and double-check that you have correct the derivative of the given function. Then set the deriv. = to 0 and solve for x.

OpenStudy (mathmale):

I need to get off the 'Net. Good luck!